# A Cost-Efficient FPGA Implementation of Tiny Transformer Model using Neural ODE

### v1

@@ 6.2.3 Effects of quantization
To reduce the memory footprint and implement the proposed model on modest-sized FPGAs, we want to apply quantization to as many layers as possible. On the other hand, since there is a trade-off between the degree of quantization and accuracy, there is a risk of significant accuracy degradation if quantization is applied to all layers without careful consideration. Here, we conduct ablation studies to investigate the impact of learning rates on the quantized blocks. We evaluate different combinations of quantized blocks. It should be noted that quantization is always applied to DSBlocks that are used only once in a single inference. We expect quantization on these blocks has less impact on accuracy. DSC is applied to ODEBlocks only. The initial and minimum learning rates for the cosine annealing scheduler are set to (5e-3, 1e-5), (5e-4, 1e-5), and (1e-4, 1e-5).
Fig. 7 shows the evaluation results. In these graphs, “DS only” indicates that the quantization is applied to DSBlocks only. “Block1”, “Block2”, and “Block3” refer to the proposed models with the quantized ODEBlock1, ODEBlock2, and MHSABlock, respectively. “Block1+2” refers to the model with both ODEBlock1 and ODEBlock2 quantized. When the learning rate is high as shown in Fig. 7(a), the training process fails to converge, leading to a significant accuracy loss especially after 150 epochs due to warm restart. In Fig. 7(b), on the other hand, the training succeeds in all the settings. In Fig. 7(c), while the accuracy improves with the number of epochs, the final accuracy of the models with quantized ODEBlock/MHSABlock is slightly lower compared to that in Fig. 7(b), suggesting that the learning rate used is not optimal. Considering that Block3 gives almost the same performance as the DS only setting and achieves higher accuracy than Block1 and Block2, MHSABlock is more tolerant to quantization than ODEBlocks. In addition, while ODEBlock2 contains more parameters than ODEBlock1, Block2 obtains better results than Block1. This implies that the having more parameters in a block does not necessarily increase its sensitivity to quantization.
(a) $lr$ = [5e-3,1e-5]
(b) $lr$ = [5e-4, 1e-5]
(c) $lr$ = [1e-4, 1e-5]
Figure 7: Ablation study on quantization and learning rate
@@ 6.2.4 FPGA resource utilization
Table 4 presents the FPGA resource utilization of the proposed model with 4-bit and 8-bit quantization applied to all building blocks. Thanks to the quantization and the 14.96$\times$ parameter size reduction by Neural ODE, the feature extraction part of our model fits within the on-chip BRAM/URAM. The quantized weight parameters are mostly stored in URAM. As a result, compared to the 8-bit quantization, the URAM utilization is further reduced by 17.7% with the aggressive 4-bit quantization.
@@
Table 5: Execution time of CPU and FPGA implementations (ms)
Model Mean Max Std
CPU (8-bit quantization) 1,596 1,676 12.73
CPU (4-bit quantization) 1,590 1,640 9.27
FPGA (8-bit quantization) 130 130 0.03
FPGA (4-bit quantization) 125 125 0.03
@@ 6.2.5 Execution time
In this section, we compare the performance of our FPGA implementation with its software counterpart running on ARM Cortex-A53 CPU. The software counterpart is implemented with PyTorch. The 8-bit and 4-bit quantizations are applied to both the implementations, respectively. Table 5 presents the mean, maximum, and standard deviation of the execution time obtained by 100 runs. In the 8-bit case, the average inference time is reduced from 1,596ms to 130ms per input feature map, yielding a 12.3$\times$ speedup. Similarly, our FPGA implementation with the 4-bit quantization runs in 125ms, which is 12.8$\times$ faster than the software implementation.
@@ 6.2.6 Energy efficiency
The power consumption of the proposed FPGA implementation is evaluated based on power reports from Xilinx Vivado. For both the 8-bit and 4-bit quantized models, the power consumption in the PS part is 2.64W, while that of the proposed IP core is 1.02W. Considering that the inference times of the 8-bit and 4-bit quantized models are accelerated by 12.3$\times$ and 12.8$\times$, the energy efficiency is improved by 8.85$\times$ and 9.21$\times$, respectively.
@@ 7 Summary
Transformers, featuring the highly expressive MHSA mechanism, have demonstrated remarkable accuracy in the field of image processing among recent AI models. This paper focuses on a hybrid model that combines Transformer and CNN architectures. Considering the resource constraints of edge devices, we employed Neural ODE, an approximation technique inspired by ResNet, to design a tiny CNN-Transformer hybrid model with 94.6% less parameters than ResNet50. The parameter size of the proposed model is the smallest among the state-of-the-art lightweight models while it offers a comparable accuracy. We proposed the resource-efficient FPGA design of the tiny Transformer model tailored for edge devices. We employed a recently-proposed LUT-based quantization technique and conducted the ablation studies to investigate the impact of quantization on different parts of the proposed model. We implemented the entire feature extraction part on a modest-sized FPGA to accelerate inference speed. Compared to an embedded CPU, our FPGA implementation achieved 12.8$\times$ faster runtime and 9.21$\times$ higher energy efficiency. Thanks to the small parameter size and quantization, the parameters fit within the limited on-chip memory, eliminating unnecessary data transfer overhead. Using high-end FPGAs would allow more design choices and further optimizations, potentially leading to greater improvements on the speed and accuracy.

### v2

@@ 6.2.3 Effects of quantization
To reduce the memory footprint and implement the proposed model on modest-sized FPGAs, we want to apply quantization to as many layers as possible. On the other hand, since there is a trade-off between the degree of quantization and accuracy, there is a risk of significant accuracy degradation if quantization is applied to all layers without careful consideration. Here, we conduct ablation studies to investigate the impact of learning rates on the quantized blocks. Different combinations of quantized blocks are evaluated. In this experiment, 4-bit and 8-bit quantizations are applied to both weights and activations for all the target blocks. It should be noted that quantization is always applied to DSBlocks that are used only once in a single inference. Quantization on these blocks has less impact on accuracy. DSC is applied to ODEBlocks only. The initial and minimum learning rates for the cosine annealing scheduler are set to [5e-3, 1e-5], [5e-4, 1e-5], and [1e-4, 1e-5].
Table 5 shows the accuracy of each combination of quantized blocks. The accuracy before the quantization is also shown. In this table, “DS only” indicates that the quantization is applied to DSBlocks only. “Block1”, “Block2”, and “Block3” refer to the proposed models with the quantized ODEBlock1, ODEBlock2, and MHSABlock, respectively. “Block1+2” refers to the model with both ODEBlock1 and ODEBlock2 quantized. The best accuracy is achieved when the quantization is applied to DSBlocks only and the learning rate settings is [5e-3, 1e-5] for both 4-bit and 8-bit implementations. Considering that Block3 gives almost the same performance as the DS only setting and achieves higher accuracy than Block1 and Block2, MHSABlock is more tolerant to quantization than ODEBlocks. In addition, while ODEBlock2 contains more parameters than ODEBlock1, Block2 obtains better results than Block1. This implies that the having more parameters in a block does not necessarily increase its sensitivity to quantization.
@@ 6.2.4 FPGA resource utilization
Table 6 presents the FPGA resource utilization of the proposed model with 4-bit and 8-bit quantization applied to all building blocks. Thanks to the quantization and the 14.96$\times$ parameter size reduction by Neural ODE, the feature extraction part of the proposed model fits within the on-chip BRAM/URAM. The 4-bit or 8-bit quantization is applied to both weights and activations for all the target blocks. In the 4-bit and 8-bit implementations, BRAMs are preferentially consumed for weights and activations, and URAMs are utilized for the portions that exceed the BRAM capacity. Thus, BRAM utilizations of the 4-bit and 8-bit implementations are almost the same while their URAM utilizations differ. As a result, compared to the 8-bit quantization, the URAM utilization is further reduced by 17.7% with the aggressive 4-bit quantization. In the case of the Zynq UltraScale+ MPSoC ZCU104 platform, the 8-bit FPGA implementation already meets the resource constraints, and the 8-bit implementation is better than the 4-bit implementation in terms of accuracy. The number of DSP slices is the same in the 4-bit and 8-bit implementations. That is, a single DSP slice is used for a single operation regardless of the quantized bit widths. The benefit of the 4-bit implementation over the 8-bit one is the parameter size reduction.
@@
Table 7: Execution time (ms) and GOPS of CPU and FPGA implementations
Model Mean Max Std GOPS
CPU (8-bit quant.) PS 1,596 1,676 12.73 0.13
CPU (4-bit quant.) PS 1,590 1,640 9.27 0.13
FPGA (8-bit quant.) PS+PL 162.1 167.5 14.40 1.30
PS 118.7 124.5 0.76 0.19
PL 43.44 43.45 0.02 4.33
FPGA (4-bit quant.) PS+PL 161.8 170.3 18.60 1.30
PS 118.6 127.6 1.06 0.19
PL 43.33 43.34 0.03 4.34
6.2.5 FPGA inference accuracy
Here, our 4-bit and 8-bit FPGA implementations are executed on the ZCU104 platform and evaluated in terms of inference accuracy. As shown in Table 5, our simulation results show that the Block3 model where the quantization is applied to DSBlocks and MHSABlock with learning rate of [5e-3, 1e-5] strikes a good balance between the accuracy and parameter size. As a result, 4-bit and 8-bit versions of the Block3 model are implemented on the ZCU104 platform as mentioned in Sec. 5.3. Please note that float32 is used for blocks other than DSBlocks and MHSABlock in the software simulation while a fixed-point number format is used for the other blocks in the FPGA implementation. To see this subtle difference, the simulation and on-board results of the 4-bit and 8-bit implementations are compared in terms of inference accuracy. The results in Table 8 demonstrate that the accuracy loss of the FPGA implementation is quite small in both the 4-bit and 8-bit cases.
Table 8: Top-1 accuracy in software and FPGA board
Platform Model 8-bit acc (%) 4-bit acc (%)
Simulation Block3 81.56 80.78
FPGA board Block3 81.53 79.18
@@ 6.2.6 Execution time
In this section, the performance of the FPGA implementation is compared with its software counterpart running on ARM Cortex-A53 CPU. The software counterpart is implemented with PyTorch. The 8-bit and 4-bit quantizations are applied to both the implementations, respectively. Table 7 presents the mean, maximum, and standard deviation of the execution time obtained by 100 runs. Please note that, in the FPGA implementation, only the pre- and post-processing is executed on the PS part while the feature extraction part is executed on the PL part. The total execution times (PS + PL) of the 8-bit and 4-bit FPGA implementations are 162.1ms and 161.8ms; thus, they achieve 9.85$\times$ and 9.83$\times$ speedups, respectively. If we look at only the feature extraction part, they achieve 34.01$\times$ and 33.96$\times$ speedups, respectively.
Table 7 also presents GOPS (giga operations per second) values based on the mean execution times. The peak performance is 41.2 GOPS when 206 DSP slices are fully-utilized for multiply and add operations at 200MHz, while the measured GOPS value of the PL part is 4.33 GOPS. The proposed accelerator is implemented as a stream of separated blocks, such as ODEBlock1, DSBlock1, ODEBlock2, DSBlock2, and MHSABlock, as shown in Fig. 6. Although our current accelerator design is modular, meaning that these blocks can be added or removed depending on the model, it would be possible to bring the measured GOPS value closer to the peak GOPS value by reducing the number of DSP slices by reusing the same DSP slices for all the blocks. Such architectural exploration and optimization are our future work.
@@ 6.2.7 Energy efficiency
The power consumption of the proposed FPGA implementation is evaluated based on power reports from Xilinx Vivado. For both the 8-bit and 4-bit quantized models, the power consumption in the PS part is 2.64W, while that of the proposed IP core is 1.02W. Considering that the inference times of the 8-bit and 4-bit quantized models are accelerated by 9.85$\times$, the energy efficiency is improved by 7.10$\times$.
@@ 7 Summary
Transformers, featuring the highly expressive MHSA mechanism, have demonstrated remarkable accuracy in the field of image processing among recent AI models. This paper focuses on a hybrid model that combines Transformer and CNN architectures. Considering the resource constraints of edge devices, we employed Neural ODE, an approximation technique inspired by ResNet, to design a tiny CNN-Transformer hybrid model with 95.1% less parameters than ResNet50. The parameter size of the proposed model is the smallest among the state-of-the-art lightweight models while it offers a comparable accuracy. We proposed the resource-efficient FPGA design of the tiny Transformer model tailored for edge devices. We employed a recently-proposed LUT-based quantization technique and conducted the ablation studies to investigate the impact of quantization on different parts of the proposed model. The entire feature extraction part is implemented on a modest-sized FPGA to accelerate inference speed. Compared to an embedded CPU, the FPGA implementation achieved a 34.01$\times$ faster runtime without pre- and post-processing and a 9.85$\times$ faster runtime for the overall inference, resulting in a 7.10$\times$ higher energy efficiency. Thanks to the small parameter size and quantization, the parameters fit within the limited on-chip memory, eliminating unnecessary data transfer overhead. Using high-end FPGAs would allow more design choices and further optimizations, potentially leading to greater improvements on the speed and accuracy. While basic forms of the attention mechanism and Neural ODE are employed in this paper, it is possible to further refine the proposed model in terms of computational efficiency and accuracy, as discussed in Sec 2. In this paper, the proposed model was implemented as a sequence of multiple blocks in the PL part. As an alternative approach, it can be implemented as a single configurable block. Although the single block implementation is more complex and challenging, it would be beneficial from a resource efficiency viewpoint, enabling a performance improvement in cooperation with more aggressive parallelization. A demonstration video of the proposed FPGA implementation on ZCU104 board is available on YouTube  [^2] https://youtu.be/jldJGnSYfVQ?si=Ss0PWq2i17hB2Q6S.